# BARE: Combining Base and Instruction-Tuned Language Models for Better Synthetic Data Generation

### v2

@@ Abstract
As the demand for high-quality data in model training grows, researchers and developers are increasingly generating synthetic data to tune and train LLMs. A common assumption about synthetic data is that sampling from instruct-tuned models is sufficient; however, these models struggle to produce diverse outputs—a key requirement for generalization. Despite various prompting methods, in this work we show that achieving meaningful diversity from instruct-tuned models remains challenging. In contrast, we find base models without post-training exhibit greater diversity, but are less capable at instruction following and hence of lower quality. Leveraging this insight, we propose Base-Refine (BARE), a synthetic data generation method that combines the diversity of base models with the quality of instruct-tuned models through a two-stage process. With minimal few-shot examples and curation, BARE generates diverse and high-quality datasets, improving downstream task performance. We show that fine-tuning with as few as 1,000 BARE-generated samples can reach performance comparable to the best similarly sized models on LiveCodeBench tasks. Furthermore, fine-tuning with BARE-generated data achieves a 101% improvement over instruct-only data on GSM8K and a 18.4% improvement over SOTA methods on RAFT.
Machine Learning, ICML
Figure 1: Histogram of pairwise embedding similarities for 1000 Llama-3.1-70B-Base vs Instruct generations of grade school math problems. The base distribution is further to the left, indicating lower similarity and hence higher diversity. Note that frequency is of the cosine similarity score.
Figure 2: Accuracy of a Llama-3.2-1B-Instruct model fine-tuned on 4 different sets of synthetically generated math problems, evaluated on a randomly selected $n=500$ subset of GSM8K. Training with BARE-generated data outperforms both Base and Instruct generations.
@@ 1 Introduction
As Large Language Models (LLMs) grow in size and capability, the demand for high-quality, diverse data in model training is outpacing human-generated data, necessitating the use of synthetically generated data (Villalobos et al., 2024). Consequently, we expect the compute spent on curating and generating data for model training to increase significantly in the coming years, especially in low-data domains where high-quality data is scarce. For example, it is common to use LLMs to generate synthetic data for a variety of tasks such as math, code, function calling, general reasoning, etc. (Yu et al., 2024; Guo et al., 2024; Patil et al., 2023; Liu et al., 2024).
The ease of data generation has led many dataset creators (Samvelyan et al., 2024; Zhang et al., 2024c; NovaSky Team, 2025) and model creators (Dubey et al., 2024; Qwen et al., 2025; Nvidia et al., 2024; Guan et al., 2025; NovaSky Team, 2025) to turn to synthetic data sampled from instruct-tuned language models as a replacement for tasks where human-generated data is lacking.
Challenges with Instruct-Tuned Models.
Though instruct-tuned models have become more and more capable at complex tasks resulting in higher quality generations, there have been discussions of how accepted methods for post-training can lead to mode collapse (Shumailov et al., 2024; Wong et al., 2024; Lambert et al., 2024). Mode collapse refers to an inability for models to generate diverse outputs to queries that don’t have a single response, yet synthetic data is primarily useful when it is both high quality and diverse (Chen et al., 2024; Raventós et al., 2023).
Intuitively, one might expect a variety of prompting techniques that explicitly encourage diversity to address this gap. For example, prompting methods we baseline against include conditioning a generation on all previous ones and prompting for it to be distinct, assuming different personas for different generations, or asking for multiple distinct entries in a single call (Zhang et al., 2024b; Naik et al., 2023; Fröhling et al., 2024). However, we find that the amount of meaningful diversity from such methods is limited and cannot fully address the gap introduced by post-training. On an open-ended generation task like email generation, such prompting methods do not meaningfully improve dataset diversity (Table 2) or trained model accuracy (Figure 9).
Base vs. Instruct-Tuned Models.
An alternative approach is to leverage base models, which are not constrained by post-training biases and thus produce outputs that better represent diversity present in real-world data (OpenAI, 2024b). More quantitatively, as shown in Figure 2, base models generate outputs with noticeably lower pairwise cosine similarity (mean similarity: $0.313$) compared to instruct-tuned models (mean similarity: $0.421$). Here we use cosine similarity as a quantitative surrogate for diversity — a dataset with greater cosine similarity has more similar items and therefore less diversity. The improved diversity in base model generations can improve downstream task performance, with Figure 2 showing that the accuracy of Llama-3.2-1B-Instruct on GSM8K increases to 22.5% from 17.8% when fine-tuning on the Llama-3.1-70B-Base generated dataset of word problems rather than the Llama-3.1-70B-Instruct generated dataset.
@@
In this work, we observe that diversity is critical for synthetic data generation when used in downstream training and fine-tuning tasks. We find that while the data generated from a base model may be more diverse, it also tends to be of lower quality (Figure 5). This can counteract the gains from diversity and hinder downstream training. Our key insight is that by combining base and instruct-tuned models, one can improve the diversity of a dataset while controlling the quality of individual data entries.
To this end, we introduce Base-Refine (BARE) — a novel approach that leverages the diverse data generated by the base models and refines it with instruct-tuned models. In a variety of settings, we find this two-stage process enhances diversity without compromising quality, enabling the generation of datasets that improve downstream performance.
We evaluate fine-tuning with BARE-generated data on contemporary tasks, such as the recently introduced Retrieval-Augmented Fine-Tuning (RAFT) method (Zhang et al., 2024c). RAFT generates synthetic question/answer pairs starting from just a corpus to fine-tune a RAG model, but we find that the diversity of data it generates is low. By replacing the data generator in RAFT with BARE, we show we can achieve up to a 18.4% fine-tuned accuracy lift over the existing work that solely uses instruct-tuned models.
As another example, BARE improves the generation of synthetic math training data. We generated grade school math problems similar to GSM8K problems with several methods, shown in Figure 2. We find that refining Llama-3.1-70B-Base data generations with Llama-3.1-70B-Instruct improves a fine-tuned model’s accuracy on GSM8K from 22.4% (base) to 29.8% (refined). Likewise, using GPT-4o as a generator out of the box (22.4%) does not perform as well as using GPT-4o as a refiner of generations from Llama-3.1-70B-Base (35.8%).
@@
    We quantitatively investigate the quality and diversity of base and instruct-tuned models for synthetic data generation across various sampling methods to motivate better system design. In Section 3, we find that base models tend to produce more diverse responses whereas instruct-tuned models offer higher quality.
@@
    Using these insights, in Section 4, we propose Base-Refine (BARE), a practical new method for generating synthetic data. We show in Section 5 that BARE consistently improves fine-tuned model performance over various baselines, including SOTA generation methods, and with only 1000 samples can lead to fine-tuned models comparable to SOTA models of similar sizes.
@@ 2 Related Work
Synthetic data is commonly used to train state of the art models like Llama 3.3 (Dubey et al., 2024), Qwen 2.5 (Qwen et al., 2025), Nemotron 4 (Nvidia et al., 2024), and o3-mini (Guan et al., 2025). However, prior work has shown its usage poses risks such as model collapse, where iterative training on low-diversity data shifts the generation distribution toward a high-probability mean, degrading both performance and diversity (Shumailov et al., 2024; Shimabucoro et al., 2024; Guo et al., 2023). Recent work has pointed towards diversity in training data improving downstream performance, though doesn’t consider the quality of the data in tandem (Chen et al., 2024). In contrast, the BARE pipeline is designed with both of these objectives in mind, producing diverse high-quality data to support model training.
@@
Current methods to improve LLM generation diversity include prompting, sampling, and multi-stage generation.
Prompting methods leverage instruction-following by requesting distinct responses in single calls or varying prompts (e.g., assigning distinct ‘personalities’ or randomizing few-shot examples) (Zhang et al., 2024b; Naik et al., 2023; Fröhling et al., 2024; Chen et al., 2024; Li et al., 2022). However, compared to BARE, these methods depend on pre-existing data or curated prompts, limiting scalability.
Sampling methods like temperature scaling and nucleus sampling (Holtzman et al., 2020) are widely used, but often prioritize token-level randomness over semantic diversity. Indeed, methods such as logit suppression (Chung et al., 2023) can enhance diversity but may require significant manual refinement to maintain quality.
Multi-Stage generation methods vary in approach. SimpleStrat (Wong et al., 2024) explicitly generates strata of the solution space and then conditions responses on each, effective for short-answer responses (often a single word or phrase), though limiting applicability to broader synthetic data generation. More complex frameworks require extensive human setup or curation to seed the diversity with topics, making them less scalable (Lambert et al., 2024; Li et al., 2024a, 2023). In contrast, BARE, is a practical two-step pipeline that we show works generally with minimal human involvement.
Prior work has studied differences in base and instruct models in calibration (OpenAI, 2024b) and agentic environments (Li et al., 2024b), though no work to our knowledge has leveraged base models for synthetic data generation.
@@
Synthetic data utility is typically assessed by downstream performance, though to motivate better systems, we additionally study diversity and entry-wise quality.
Token-level metrics like self-BLEU (Zhu et al., 2018) are common, though embedding-based approaches (e.g., BERTScore (Zhang et al., 2019), Sentence-BERT (Reimers and Gurevych, 2019)) better capture semantic diversity rather than token diversity, which is our primary focus.
Lastly, while dataset-wide quality is often measured via downstream performance, assessing individual synthetic samples remains underexplored. In Section 3, we introduce an entry-wise quality measure to evaluate sample realism, ensuring robust synthetic data generation.
@@ 3 Motivation
Synthetic data generation should result in a diverse dataset with high-quality entries. Thus we motivate our design of BARE with investigations on the diversity and entry-wise quality of synthetically generated data from base and instruct-tuned models, independent of downstream utility.
3.1 Diversity & Quality
@@
Following Tevet and Berant (2021), we use the average neural similarity score to measure the diversity of a generated dataset. Specifically, we use OpenAI’s text-embedding-3-small (OpenAI, 2024a) to generate embeddings and use cosine similarity to calculate similarity scores, as recommended by OpenAI. We calculate pairwise cosine similarity scores for items in a generated dataset and analyze the resulting distribution of similarities. A lower average similarity indicates a more diverse dataset.
@@
We seek a metric measuring the quality of individual entries in the generated datasets. To this end, we propose the indistinguishability rate (IR) of the entries as a quality metric. We believe that we are the first to propose this metric for the quality of an individual data entry. Specifically, we measure how often a strong LLM (e.g. GPT-4o) fails to identify the synthetically generated entry as being of lower quality when combined with $n=3$ real dataset entries. A high IR indicates that generated data closely matches properties of real-world data and the discriminator is randomly guessing, while a low IR indicates the generated data are out-of-distribution from real world data. An example IR prompt is available in Appendix B.
@@ 3.2 Experimental Setup
To investigate diversity and quality differences between instruct-tuned and base models, we evaluate a synthetic dataset generated using Llama-3.1-70B-Instruct and Llama-3.1-70B-Base (Dubey et al., 2024). We additionally use GPT-4o (OpenAI, 2024c) as an instruct-tuned model with strong instruction-following capabilities.
To allow fair comparisons and simulate a low-data setting we always provide three few-shot examples.
@@
We examine a variety of domains covering many tasks. Our data generation domains are:
@@
  • 20 Newsgroups (Pedregosa et al., 2011) generating training data for classifying Usenet messages into one of 20 newsgroup sources. The model generates classes along with content, allowing the generation method to determine the class distribution of the synthetic dataset.
  • Retrieval-Augmented Fine-Tuning (RAFT) (Zhang et al., 2024c), a domain and generator-agnostic synthetic data generation framework for fine-tuning data in RAG tasks. Here, Q/A pair data generation needs to be conditioned on contexts mimicking retrieval results. We use:
@@
  • GSM8K (Cobbe et al., 2021), generating grade-school math word problems and solutions for fine-tuning.
@@
For the classification tasks of Enron and Newsgroups, we generate a dataset of size $n=500$. For the generative model fine-tuning tasks of HotpotQA, PubMedQA, GSM8K, and LCB TOP, we generate a larger dataset of size $n=1000$.
We sample at a default temperature of $0.7$ for base models and the highest temperature at which we experimentally found data generation is still coherent for instruct models, which is $1.0$ for Llama models and $1.2$ for GPT-4o. However, for Enron, we sample from GPT-4o at a temperature of $1.0$ to maintain generation coherence.
Note we do not measure diversity in the RAFT domains as we expect noise due to conditioning on different documents/contexts. For Enron, due to the relatively large differences between spam and legitimate emails, only similarities between emails of the same type are calculated.

### v3

# BARE: Leveraging Base Language Models for Few-Shot Synthetic Data Generation

@@ Abstract
As the demand for high-quality data in model training grows, researchers and developers are increasingly generating synthetic data to tune and train LLMs. However, current data generation methods rely on seed sets containing tens of thousands of examples to prompt instruction-tuned models. This reliance can be especially problematic when the curation of high-quality examples is expensive or difficult. In this paper we explore the novel few-shot synthetic data generation setting – generating a high-quality dataset from a few examples. We show that when working with only a few seed examples, instruction-tuned models used in current synthetic data methods produce insufficient diversity for downstream tasks. In contrast, we show that base models without post-training, largely untapped for synthetic data generation, offer substantially greater output diversity, albeit with lower instruction following abilities. Leveraging this insight, we propose Base-Refine (BARE), a novel two-stage method that combines the diversity of base models with the quality assurance of instruction-tuned models. BARE excels in few-shot synthetic data generation: using only 3 seed examples it generates diverse, high-quality datasets that significantly improve downstream task performance. We show that fine-tuning Llama 3.1 8B with 1,000 BARE-generated samples achieves performance comparable to state-of-the-art similarly sized models on LiveCodeBench tasks. Furthermore, data generated with BARE enables a 101% improvement for a fine-tuned Llama 3.2 1B on GSM8K over data generated by only instruction-models, and an 18.4% improvement for a fine-tuned Llama 3.1 8B over the state-of-the-art RAFT method for RAG data generation.
@@ 1 Introduction
As Large Language Models (LLMs) grow in size and capability, the demand for high-quality, diverse data in model training is outpacing human-generated data, necessitating the use of synthetically generated data (Villalobos et al., 2024). Already, it is common to use LLMs to generate synthetic data for a variety of tasks such as math, code, function calling, and general reasoning (Yu et al., 2024; Guo et al., 2024; Patil et al., 2023; Liu et al., 2024). The convenience of data generation has led many dataset creators (Samvelyan et al., 2024; Zhang et al., 2024b; NovaSky Team, 2025) and model creators (Dubey et al., 2024; Qwen, 2025; Nvidia, 2024; Guan et al., 2025; NovaSky Team, 2025) to turn to synthetic data sampled from instruction-tuned language models as a replacement for tasks where human-generated data is lacking.
Figure 1: Histogram of pairwise embedding cosine similarity scores for 1000 Llama-3.1-70B-Base vs Instruct generations of grade school math problems with 3 seed examples. The base distribution is further left, indicating lower similarity and hence higher diversity.
Figure 2: Accuracy of a Llama-3.2-1B-Instruct model fine-tuned on real-world data and 5 different sets of math problems synthetically generated using 3 seed examples, evaluated on a randomly selected $n=500$ subset of GSM8K. Training with BARE-generated data outperforms all other data sources.
Synthetic data is most valuable when it is both high quality and diverse (Chen et al., 2024; Raventós et al., 2023). However, while instruction-tuned models excel at generating high-quality outputs, they struggle with generating diverse outputs. Recent work suggests that this lack of diversity is a consequence of post-training, where standard techniques lead to mode collapse (Shumailov et al., 2024; Wong et al., 2024; Lambert et al., 2024), limiting a model’s ability to generate varied responses to open-ended queries.
To achieve sufficient diversity with instruction-tuned models, prevailing methods require developers to provide large and diverse seed sets. For instance, OSS-Instruct (Wei et al., 2024) uses 80,000 code snippets and Humpback (Li et al., 2024c) over 500,000 text segments. However, this reliance imposes significant burdens on developers through the need for data curation, ultimately hindering adoption. While this inefficiency is a general challenge, it becomes especially acute where large seed sets are difficult to obtain (e.g., specialized tasks such as clinical research (Sivarajkumar et al., 2023)). Additionally, we show that methods to induce diversity via prompting methods (e.g. including past examples in context) (Zhang et al., 2024a; Naik et al., 2023; Fröhling et al., 2024) are insufficient for few-shot generation. Thus, a critical need exists for methods that can generate high-quality, diverse data in few-shot settings (i.e., with minimal seed input).
Base Models as an Alternative.
Given the limitations of instruction-tuned models in few-shot scenarios, an underexplored alternative source of diversity is the use of base models. Not being subject to the same post-training procedures, base models may produce outputs that better reflect the diversity found in real-world data OpenAI (2024b). Quantitatively, base models demonstrate greater diversity. As shown in Figure 2, they generate outputs with noticeably lower pairwise embedding cosine similarity (mean 0.313) compared to instruction-tuned models (mean 0.421); lower cosine similarity indicates greater diversity. This increased diversity can benefit downstream tasks: Figure 2 shows that fine-tuning Llama-3.2-1B-Instruct on GSM8K using data from Llama-3.1-70B-Base yields a 22.5% accuracy, compared to 17.8% using data from Llama-3.1-70B-Instruct.
However, this advantage comes with a trade-off. Due to their weaker instruction-following capabilities, base model generations often suffer from lower quality (Figure 4), which can negate the benefits of increased diversity and hinder downstream training performance. Our key insight is that by managing the quality of individual data points using instruction-tuned models, we can harness the diversity advantage of base models effectively.
@@
Leveraging the distinct strengths of base and instruction-tuned models, we introduce Base-Refine (BARE): a novel approach for few-shot synthetic data generation. BARE leverages a base model to generate diverse initial generations with minimal seed data and refines each initial generation according to desiderata with instruction-tuned models. In a variety of few-shot settings, we show our two-stage process enhances diversity without compromising quality, enabling the generation of datasets that improve downstream performance using only 3 seed examples.
On the LiveCodeBench Test Output Prediction task, fine-tuning with 1,000 BARE-generated examples achieves performance comparable to state-of-the-art models of similar size. Further, when applied to the Retrieval-Augmented Fine-Tuning (RAFT) method (Zhang et al., 2024b)—a state-of-the-art technique for generating synthetic Q&A pairs over retrieved documents for RAG—replacing RAFT’s instruction-tuned only generator with BARE improves the fine-tuned model’s accuracy by up to 18.4% over the original RAFT approach (Figure 8). In mathematical reasoning, as exemplified by synthetic GSM8K-style problems, BARE also demonstrates considerable benefits. Using Llama-3.1-70B-Base for initial generation and GPT-4o for refinement with BARE increases a fine-tuned Llama-3.2-1B-Instruct model’s accuracy on GSM8K to 35.8%. Training on BARE-generated data significantly surpasses the 22.4% accuracy achieved with instruct-only generation using GPT-4o. Furthermore, it exceeds the 26.6% accuracy obtained by training on examples from the official human-generated GSM8K training set (Figure 2).
@@
    We quantitatively investigate the quality and diversity of base and instruction-tuned models for synthetic data generation across various sampling methods to motivate better system design. We show that base models tend to produce far more diverse responses whereas instruction-tuned models offer higher quality.
@@
    Building on these insights, we propose Base-Refine (BARE), a practical method novelly leveraging base models for the under-explored task of few-shot synthetic data generation. We demonstrate that BARE consistently outperforms past methods on small seed sets, including SOTA generation methods. Generating from as few as 3 seed examples, BARE can yield fine-tuned models comparable to SOTA models of similar sizes trained with much larger seed sets.
@@ 2 Related Work
Synthetic data is commonly used to train state of the art models like Llama 3.3 (Dubey et al., 2024), Qwen 2.5 (Qwen, 2025), Nemotron 4 (Nvidia, 2024), and o3-mini (Guan et al., 2025). However, prior work has shown its usage poses risks such as model collapse, where iterative training on low-diversity data shifts the generation distribution toward a high-probability mean, degrading both performance and diversity (Shumailov et al., 2024; Shimabucoro et al., 2024; Guo et al., 2023). Recent work indicates that diversity in training data can improve downstream performance (Chen et al., 2024). However, this research often does not simultaneously consider the quality of the data in tandem. In contrast, the BARE pipeline is designed with the twin objectives of diversity and quality in mind, producing diverse, high-quality data to support model training.
@@
Sampling methods like temperature scaling and nucleus sampling (Holtzman et al., 2020) are widely used to improve diversity, but often prioritize token-level randomness over semantic diversity. Indeed, methods such as logit suppression (Chung et al., 2023) can enhance diversity but may require significant manual refinement to maintain quality unlike BARE which maintains quality as a first order requirement.
Many current synthetic data generation processes instead rely on varying the prompt to elicit a diverse dataset. One common approach is to rely on heavily curated prompts limiting scalability (Zhang et al., 2024a; Naik et al., 2023; Fröhling et al., 2024; Chen et al., 2024; Li et al., 2022). Another effective approach is to utilize a large diverse seed set, up to 100,000s of examples in size (Lambert et al., 2024; Li et al., 2023; Wei et al., 2024; Li et al., 2024c), requiring significant curation effort. BARE obviates both of these concerns by focusing on the few-shot setting (we use only 3 seed examples) and showing significant performance gains without prompt engineering.
Others have explored using an LLM to generate the seed set (Wong et al., 2024; Li et al., 2024a), but still use instruction-tuned models. These methods thus still suffer from the limited diversity of instruction-tuned models. In contrast, BARE utilizes base models to generate a large and diverse seed set. To our knowledge, no prior work has focused on this approach. Though prior work has studied differences in base and instruct models in calibration (OpenAI, 2024b) and agentic environments (Li et al., 2024b), no work has leveraged base models for synthetic data.
@@
The utility of synthetic data is typically assessed by downstream performance. However, to motivate the development of better systems, we also study diversity and the quality of individual entries (entry-wise quality). Token-level metrics like self-BLEU (Zhu et al., 2018) are common, though embedding-based approaches that are commonly used (e.g., BERTScore (Zhang et al., 2019), Sentence-BERT (Reimers and Gurevych, 2019)) better capture semantic diversity rather than token diversity, which is our primary focus. Lastly, while dataset-wide quality is often measured via downstream performance, assessing individual synthetic samples remains underexplored. In Section 3, we introduce an entry-wise quality measure to evaluate sample realism, ensuring robust synthetic data generation.
@@ 3 Motivation
Synthetic data generation should result in a diverse and high-quality dataset. In this section, we run experiments to demonstrate the challenges of few-shot synthetic data generation with current methods and the differences between base and instruction tuned models. We demonstrate the potential of base models as a source of diversity, motivating the design of BARE.
3.1 Diversity & Quality Metrics
@@
Following Tevet and Berant (2021); Cann et al. (2023); Cox et al. (2021), we use the average neural similarity score to measure the diversity of a generated dataset. Specifically, we use OpenAI’s text-embedding-3-small (OpenAI, 2024a) to generate embeddings and use cosine similarity to calculate similarity scores, as recommended by OpenAI. We calculate pairwise cosine similarity scores for items in a generated dataset and analyze the resulting distribution of similarities. A lower average similarity indicates a more diverse dataset.
@@
To measure entry-wise quality, we propose the indistinguishability rate (IR), a novel metric inspired by the adversarial framework of Generative Adversarial Networks (GANs) (Goodfellow et al., 2014). We task a strong LLM (e.g., GPT-4o) as a "discriminator" to distinguish a synthetic entry from $n=3$ real entries. The IR is the rate at which it fails to distinguish the synthetic entry. A high IR suggests the synthetic data closely mimics real data, while a low IR indicates it is easily identifiable as out-of-distribution. An example IR prompt is in Appendix D.
@@ 3.2 Experimental Setup
To investigate diversity and quality differences between instruction-tuned and base models, we evaluate synthetic datasets generated using Llama-3.1-70B-Instruct and Llama-3.1-70B-Base (Dubey et al., 2024). We also use GPT-4o (OpenAI, 2024c) as a stronger instruction-tuned model.
@@
To analyze the few-shot setting, each generation method is exposed to only three real-world examples from the following established benchmarks:
@@
  • 20 Newsgroups (Pedregosa et al., 2011) generating training data for classifying Usenet messages into one of 20 newsgroup sources. We generate classes along with content, allowing the generation method to determine the class distribution of the synthetic dataset.
  • Retrieval-Augmented Fine-Tuning (RAFT) (Zhang et al., 2024b), a domain and generator-agnostic synthetic data generation framework for fine-tuning data in RAG tasks. Q/A pair generation is conditioned on contexts mimicking retrieval results from a corpus. We use:
@@
  • GSM8K (Cobbe et al., 2021), generating grade-school math problems and solutions for fine-tuning.
@@
For the classification tasks of Enron and Newsgroups, we generate a dataset of size $n=500$. For the generative model fine-tuning tasks of HotpotQA, PubMedQA, GSM8K, and LCB TOP, we generate a larger dataset of size $n=1000$. For the RAFT domains, this would mean 10 questions for each of 100 simulated retrievals; we do not compute diversity for RAFT due to differences in retrievals.
We use straight-forward prompts with limited domain-tailoring (examples in Appendix D). Note that our prompt for instruction-tuned models is comparable to those used in current synthetic data generation techniques such as OSS-Instruct (Wei et al., 2024). Indeed, validation training runs adapting OSS-Instruct’s prompts to our setting yield comparable results to our prompts. We thus use our instruction-tuned-only methods as a stand-in for existing methods. Further sampling details, such as temperature, are in Appendix A.
Figure 3: Distribution of pairwise embedding cosine similarity scores for Llama-3.1-70B-Instruct and Llama-3.1-70B-Base generations for Enron spam. Base model distribution have more density in the low-similarity region and less density in the high-similarity region, indicating greater diversity.